OpenStudy (iceman35):

Meg and her sister are going to make 30 sandwiches for the family picnic. If Meg can make a sandwich in 35 seconds and her sister can make a sandwich in 29 seconds how long would it take both girls to complete tha job?

OpenStudy (iceman35):

@pooja195 HEY POOJA HALP ME XD

OpenStudy (phi):

I use the idea of rate*time = amount any idea what the rate of Meg is ?

OpenStudy (iceman35):

Meg is one sandwich per 35 seconds right?

OpenStudy (phi):

yes, 1/35 and 1/29 for the other if they don't slow each other down, we can add their rates \[ \left(\frac{1}{35}+\frac{1}{29} \right)t= 30 \]

OpenStudy (iceman35):

Oh ok so I add those and divide by 30 to get the answer

OpenStudy (phi):

you would get a number A (call it) A t = 30 divide both sides by A t = 30/A

OpenStudy (iceman35):

D: I don't get it what does A stand for

OpenStudy (phi):

just a number (i.e. the sum of 1/35 + 1/29) the idea is you divide both sides of the equation by that number to "solve for t"

OpenStudy (iceman35):

oh ok what do I do after that step?

OpenStudy (phi):

what did you get so far ?

OpenStudy (iceman35):

Thank you for your help but im going to be completely honest and say idk whats going on

OpenStudy (phi):

1/35+1/29 = 64/1015

OpenStudy (iceman35):

oh right so t=30/A A=64/1015 t=30*1015/64= 3045/64

OpenStudy (phi):

\[ \left(\frac{1}{35}+\frac{1}{29} \right)t= 30 \\ \frac{65}{1015}t= 30 \]

OpenStudy (phi):

you mean 30450/64 or 475.78 seconds

OpenStudy (phi):

that is about 8 minutes

OpenStudy (iceman35):

hmm Thank you so much :D:D I think I can apply your thoughts onto my next homework problem